27 125 343 1331 ? 4913
Answer by stanbon(75887)

27 125 343 1331 ? 4913
-----------
3^3 = 27
5^3 = 125
7^3 = 343
11^3 = 1331
? = 13^3 = 2197
17^3 = 4913
----------
Cheers,
Stan H.
------------